OpenStudy (daggiemib):

Roast beef has 25g of protein and 11g of calcium per serving. A serving of mashed potatoes has 2g of protein and 25g of calcium. How many servings of each are needed to supply exactly 29g of protein and 61g of calcium?

OpenStudy (trojanpoem):

Ok, turn it into algebraic equations: 25x + 2y = 29 ( 25 g of protein * the number of servings + 2 * number of servings = total protein) 11x + 25y = 61 Solve both equations together and you have number of servings , do you know how to solve them ?

OpenStudy (daggiemib):

what about calcium

OpenStudy (daggiemib):

i know how to do linear systems im just not sure how they should be set up

OpenStudy (trojanpoem):

The second equations Roast beef gives 11g each serving mashed potatos gives 25g each serving so 11 * number of roast beef servings + 25 * number of mashed potatos servings = total calcuim

OpenStudy (daggiemib):

oh so ill do two sets ? and then substitute ?

OpenStudy (trojanpoem):

25x + 2y = 29 11x + 25y = 61 You could use elimination , or substitution: Multiply first equation * -11 ( negative of second x coefficient) and second equation times 25 -275x - 22y = -319 275x +625y = 1525 603y = 1206 y = 2 11x + 50 = 61 11x = 11 x = 1

OpenStudy (daggiemib):

ohhhh okay thank you so much

OpenStudy (trojanpoem):

Any time.